# Privacy-Aware Semantic Cache for Large Language Models

### v1

@@ 4.4. MeanCache Embedding Compression utility and Impact on Storage Space
Figure 8 illustrates the effects of MeanCache dimension reduction utility on the storage, semantic matching speed (overhead), and MeanCache’s performance (F score). The x-axis indicates the number of queries stored in the cache, while the y-axis shows storage size, average search time, and the F score in respective graphs. MeanCache-Compressed (MPNet) and MeanCache-Compressed (Albert) represent instances where MeanCache decreases the embedding dimensions from 768 to 64 by employing the compression, as detailed in MeanCache design (Section 3.2).
Figure 8 demonstrates that an increase in stored queries linearly raises storage needs. Yet MeanCache with compression enabled drastically lowers storage needs by 83% compared to GPTCache. Figure 8 also indicates that compression decreases the average search time, with MeanCache enabled compression being approximately 11% faster than without. Moreover, despite a slight decrease in F score with compression enabled, MeanCache still surpasses GPTCache. Furthermore, given the evidence from Figures 1 and 2 that MPNet produces more precise embeddings, and it is also clear from Figure 8 that MPNet’s embeddings are particularly resilient to compression and excel in semantic matching.
@@ 5. Related Work
In the above studies, the caching systems are designed for traditional search engines that process keyword queries (i.e., key word matching) and return a list of links to the user as a response. However, when such caching systems are applied to LLMs based web servers or APIs, they will not be able to provide a single concise response and may return many false results. On top of that, such caching techniques will fail to capture the semantic similarity among repeated queries and will result in a significantly low hit rate. For instance, if a user sends the query ”How can I increase the battery life of my smartphone?” and later sends a similar query ”Tips for extending the duration of my phone’s power source.”, the existing caches will not be able to detect the semantic similarity between these two queries and will treat them as two different queries. Zhu et al. (2023) and (Bang, 2023) propose server-side caching for the services based on LLMs to reduce the massive computational cost of LLMs. For instance, Zhu et al. (2023) upon receiving a new query at the server checks if the query is semantically similar to any query in the cache. If the query is similar to any query in the cache then the server returns the response from the cache. Otherwise, a model multiplexer is used to select the most suitable LLM for the query.
Although, these techniques can handle the semantic similarity among the queries and can provide a single concise response, they raise privacy concerns as they involve storing user queries on the server side. Second, these techniques are static and unable to adapt to the behavior of each user. Furthermore, an end user will still be charged for the query even if the query is served from the cache. Thus, we need a user-centered semantic cache that can run on the user side that can benefit the user in terms of privacy, cost, and latency. Such a cache should be able to detect the semantic similarity among the queries and should be able to adapt to the behavior of each user in a privacy preserving manner. MeanCache provides aforementioned benefits without violating user privacy.
@@ 6. Conclusion
MeanCache is the first user-centric semantic cache for Language Model (LLM) based web services, such as ChatGPT. In MeanCache, clients train the global embedding model using federated learning on their local data, while preserving user privacy. The global model formed after the aggregation generates high-quality embeddings for semantic matching. When a new user’s query is duplicated, MeanCache semantically matches the new query with the user’s local cache and returns the most relevant results. This process reduces the computational cost of the LLM service, improves bandwidth and latency, and conserves the user’s query quota for the LLM service. Despite compressing embeddings and saving 83% of storage space, MeanCache outperforms the existing baseline semantic cache (GPTCache). Therefore, MeanCache presents a promising approach for training high-quality embedding models, preserving user privacy, and enhancing the performance of LLM-based web services by serving repeated queries from the user’s local cache with a significant low rate of false hits.

### v2

@@ 5.5. MeanCache Embedding Compression utility and Impact on Storage Space
Figure 12 illustrates the effects of MeanCache dimension reduction utility on the storage, semantic matching speed (overhead), and MeanCache’s performance (F score). The x-axis indicates the number of queries stored in the cache, while the y-axis shows storage size, average search time, and the F score in respective graphs. MeanCache-Compressed (MPNet) and MeanCache-Compressed (Albert) represent instances where MeanCache decreases the embedding dimensions from 768 to 64 by employing the compression, as detailed in MeanCache design (Section 4).
Figure 12 demonstrates that an increase in stored queries linearly raises storage needs. Yet MeanCache with compression enabled drastically lowers storage needs by 83% compared to GPTCache.
Figure 12 also indicates that compression decreases the average search time, with MeanCache enabled compression being approximately 11% faster than without. Moreover, despite a slight decrease in F score with compression enabled, MeanCache still surpasses GPTCache. Furthermore, given the evidence from Figures 5 and 6 that MPNet produces more precise embeddings, and it is also clear from Figure 12 that MPNet’s embeddings are particularly resilient to compression and excel in semantic matching.
@@ 6. Related Work
In all of the above studies, the caching systems are designed for traditional search engines that process keyword queries (i.e., key word matching) and return a list of links to the user as a response. However, such caching systems, when applied to LLMs-based web servers or APIs, do not provide a single concise response and may return many false results. On top of that, such caching techniques will fail to capture the semantic similarity among repeated queries and will result in a significantly low hit rate. Zhu et al. (2023) and (Bang, 2023) propose server-side caching for the services based on LLMs to reduce the massive computational cost of LLMs. For instance, Zhu et al. (2023) upon receiving a new query at the server checks if the query is semantically similar to any query in the cache. If the query is similar to any query in the cache then the server returns the response from the cache. Otherwise, a model multiplexer is used to select the most suitable LLM for the query.
Although these techniques can handle the semantic similarity among the queries and can provide a single concise response, they raise privacy concerns as they store user queries on an external server. These techniques are static and unable to adapt to the behavior of each user. Furthermore, an end user will still be charged for the query even if the query is served from the cache. Thus, we need a user-centered semantic cache that can run on the user side, which can benefit the user in terms of privacy, cost, and latency. Such a cache should be able to detect the semantic similarity among the queries and should be able to adapt to the behavior of each user in a privacy preserving manner. MeanCache provides the aforementioned benefits without violating the user’s privacy.
@@ 7. Conclusion
MeanCache is the first user-centric semantic cache for Language Model (LLM) based web services, such as ChatGPT. In MeanCache, clients train the global embedding model using federated learning on their local data, while preserving user privacy. The global model formed after the aggregation generates high-quality embeddings for semantic matching. When a new user’s query is duplicated, MeanCache semantically matches the new query with the user’s local cache and returns the most relevant results. This process reduces the computational cost of the LLM service, improves bandwidth and latency, and conserves the user’s query quota for the LLM service. Despite compressing embeddings and saving 83% of storage space, MeanCache outperforms the existing baseline semantic cache (GPTCache). With its distribution cache design, MeanCache offers pathways to reduce one-third of the LLM query inference costs related to semantically similar queries.